OpenStudy (anonymous):

Solve this Equation y/4- 1 1/2= y/6

OpenStudy (anonymous):

y/4 - y/6 = 11/2 by rearranging terms 6y-4y / 24 = 11/2 2y/24 = 11/2 y/12 =11/2 y/6 = 11 y=66

OpenStudy (anonymous):

thats what i got but it was wrong.

OpenStudy (anonymous):

not able to find whats wrong...